John Fleming, chief administrator for Valley View Hospital, is concerned about the costs for tests in the hospital's lab. Charges for lab tests are consistently higher at Valley View than at other hospitals and have resulted in many complaints. Also, because of strict regulations on amounts reimbursed for lab tests, payments received from insurance companies and governmental units have not been high enough to cover lab costs Mr. Fleming has asked you to evaluate costs in the hospital's lab for the past month. The following information is available a Two types of tests are performed in the lab-blood tests and smears. Durina the past month, 200 blood tests and 2.000 smears were performed in the lab. b. Small glass plates are used in both types of tests. During the past month, the hospital purchased 10,000 plates at a cost of $32.900 1.000 of these piates were unused at the end of the monthi no plates were on hand at the beginning of the month. C During the past month. 1,000 hours of labor time were recorded in the lab at a cost of $10,300. d. The lab's variable overhead cost last month totaled $6.600. Valley View Hospital has never used standard icosts By searching industry iterature, however, you have determined the following nationwide averages for hospital labs: Piates Three plates are required per lao test These plates cos$t S50 each and are disposed of after the test is completed Labor Each blood test should recuire 06 hours to complete and each smear shoulc require 0.30 hours to complete. The average Cost of this lab time is $11.00 per hour Overhead:Overhead cost is basec on cirect labor-hoursThe average rate forvariable overhead is $610 per hour Peauired:
Required: 1 Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month 2. For labor cost in the lab: Compute a labor rate variance and a labor efficiency variance. b. In most hospitals, one-half of the workers in the lab are senior technicians and one-half are assistants. In an effort to reduce costs, Valley View Hospital employs only one-fourth senior technicians and three-fourths assistants. Would you recommend that this policy be continued? a. 3-a. Compute the variable overhead rate and efficiency variances 3-b. Is there any relation between the variable overhead efficiency variance and the labor efficiency variance? Complete this question by entering your answers in the tabs below. Reg 3B Reg 3A Reg 28 Rea 2A Req 1 Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month. (Indicate the effect of each variance by selecting "F for favorable, "U" for unfavorable, and "None" for no effect he zero variance) Input all amounts as positive values.) Materials price variance Materials quantity variance
month. 2. For labor cost in the lab: a. Compute a labor rate variance and a labor efficiency variance b. In most hospitals, one-half of the workers in the lab are senior technicians and one-half are assistants. In an effort to reduce costs Valley View Hospital employs only one-fourth senior technicians and three-fourths assistants. Would you recommend that this policy be continued? 3-a. Compute the variable overhead rate and efficiency variances. 3-b. Is there any relation between the variable overhead efficiency variance and the labor efficiency variance? Complete this question by entering your answers in the tabs below. Req 2B Reg 3A Req 3B Req 1 Req 2A For labor cost in the lab, Compute a labor rate variance and a labor efficiency variance. (Indicate the effect of each variance by selecting "F" for favorable, "U for unfaverable, and "None for no effect (i.e., zero variance). Input all amounts as positive vatues Labor rate variance Labor efficiency variance
Required: 1. Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month 2. For labor cost in the lab. a. Compute a labor rate variance and a labor efficiency variance. b. In most hospitals, one-half of the workers in the lab are senior technicians and one-half are assistants. In an effort to reduce costs Valley View Hospital employs only one-fourth senior technicians and three-fourths assistants. Would you recommend that this policy be continued? 3-a. Compute the variable overhead rate and efficiency variances. 3 b. Is there any relation between the variable overhead efficiency variance and the labor efficiency variance? Complete this question by entering your answers in the tabs below. Req 3A Req Req 3B Reg 2B Req 2A Compute the variable overhead rate and efficiency variances. (Indicate the effect of each variance by selecting " for favorable "U for unfavorable, and "None" for no effect (i.e, zero variance) Input all amounts as positive values.) Variable overhead rate variance Veriable overhead efficiency variance Req 38
Answer #1
1 Material price variance = AQ x (AR - SR) = 10000 x (3.29 - 3.50) = 10000 x -0.21 = $2100 F
AR = $32900/10000 = $3.29
Materials quantity variance = SR x (AQ - SQ) = 3.50 x (9000 - 6600) = 3.50 x 2400 = $8400 U
AQ = 10000 - 1000 = 9000
SQ = (200 + 2000) x 3 = 6600
2a. Labor rate variance = AH x (AR - SR) = 1000 x (10.30 - 11.00) = 1000 x -0.70 = $700 F
AR = $10300/1000 = $10.30
Labor efficiency variance = SR x (AH - SH) = 11.00 x (1000 - 720) = 11.00 x 280 = $3080 U
SH = (200 x 0.6) + (2000 x 0.3) = 120 + 600 = 720
b. No.
This could be the reason for the favorable labor rate variance but unfavorable labor efficiency variance. Hence this policy should not be continued.
3-a. Variable overhead rate variance = AH x (AR - SR) = 1000 x (6.60 - 6.10) = 1000 x 0.50 = $500 U
AR = $6600/1000 = $6.60
Variable overhead efficiency variance = SR x (AH - SH) = 6.10 x (1000 - 720) = 6.10 x 280 = $1708 U
3-b. Yes
Both are based on direct labor hours.